OpenStudy (anonymous):

(g-f)(x), f(x)=square root of x, g(x)=2x+3. I need help solving it?

OpenStudy (gorv):

(g-f)(x)=g(x)-f(x)

OpenStudy (anonymous):

How would I subtract the \[\sqrt{x}\]

Miracrown (miracrown):

well...since there are no like terms all you can do is write the difference out

OpenStudy (gorv):

leave as it is

Miracrown (miracrown):